User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Observers Meeting 23.06.2015

Current affairs:

  • Note 26m for RD1504 tonight. Don't forget to move the feed. Ho-Cd fringe check prior.
  • Hb experiments moving to Ho from late July for 2-3 months during broadband tests. Dates TBD.
  • NCRIS funding cuts and their implications
  • Correlator reports from last r1 and r4 sessions:
    1. Yg is fine
    2. Ke has unstable pcal:
      • Over several scans to all experiment
      • Amplitude is too low/null
      • Phase changes abruptly mid-scan
    3. Hb often has low fringe amplitudes in some channels that are subsequently flagged, from most frequent to least: XR5U, SR5U&SR6U, XR8L
  • windke problem? Dave had a false windstow at Ke last night — windstowed with 2km/h winds! Rebooting the wind device fixed it, but worth keeping an eye on.
  • DBBC at Ke needed to be reconfigured twice in the last few days.
  • Rxmon at Hb: Jamie's new Rxmon script is much more resilient than the old one although occasionally the connection with the fs is dropped and an alarm is sounded. This seems to be on the fs end of things however, and the connection is restored soon after without restarting Rxmon.


  • Questions?
  • Swaps?
/home/www/auscope/opswiki/data/pages/operations/agenda2015_06_23.txt · Last modified: 2015/06/24 02:59 by David Horsley